# Decoupling Layout from Glyph in Online Chinese Handwriting Generation

### v2

@@ Abstract
Text plays a crucial role in the transmission of human civilization, and teaching machines to generate online handwritten text in various styles presents an interesting and significant challenge. However, most prior work has concentrated on generating individual Chinese fonts, leaving complete text line generation largely unexplored. In this paper, we identify that text lines can naturally be divided into two components: layout and glyphs. Based on this division, we designed a text line layout generator coupled with a diffusion-based stylized font synthesizer to address this challenge hierarchically. More concretely, the layout generator performs in-context-like learning based on the text content and the provided style references to generate positions for each glyph autoregressively. Meanwhile, the font synthesizer which consists of a character embedding dictionary, a multi-scale calligraphy style encoder and a 1D U-Net based diffusion denoiser will generate each font on its position while imitating the calligraphy style extracted from the given style references. Qualitative and quantitative experiments on the CASIA-OLHWDB demonstrate that our method is capable of generating structurally correct and indistinguishable imitation samples.
@@ 1 Introduction
In this work, we aim to generate stylized and coherent online Chinese handwritten text lines based on specified text contents and given style reference samples. A handwritten line of text can be viewed as composed of individual characters and their arrangement. Based on this observation, we propose a hierarchical method that disentangles the layout generation from the writing of individual characters during the generation of full lines. We design a novel text line layout generator to arrange positions for each character based on their categories and given handwriting style references. For font generation, we construct a 1D convolutional U-Net (Ronneberger et al., 2015) network and design a multi-scale feature fusion module to fully mimic the calligraphy characteristics of the given reference sample. We utilize the publicly available CASIA-OLHWDB dataset (Liu et al., 2011) for training and evaluating our method, showcasing its effectiveness through a combination of quantitative and qualitative experiments. Figure 1 shows several visualizations of our generated samples.
Our contributions can be summarized as: 1) We propose a hierarchical approach to solve the unexplored online handwritten Chinese text line generation task. 2) We introduce a simple but effective layout generator that can generate character positions based on the text contents and writing style through in-context-like learning. 3) We construct a 1D U-Net network for font generation and design a multi-scale contrastive-learning-based style encoder to improve the ability of calligraphy style imitation.
@@ 4.2 Font Generation
Baselines and Results. There are mainly two types of online handwritten Chinese character generation approaches: the first type comprises purely data-driven methods that learn the structure of characters and calligraphy style only from the training set. The second type requires a standard printed font image as an auxiliary input, and train models to transform the standard fonts into characters with a specific writing style. The advantage of the second method is that it is easier to generate correct structures, but it cannot generate characters without corresponding standard glyphs and the computational cost will be heavier. As shown in Table 1, being a method of the first type, our model achieves state-of-the-art performance in all pure data-driven approaches and is comparable to state-of-the-art style transfer-based methods.
@@ 4.3.2 Quantitative Results
   Gaussian 0.074 0.151 0.092 0.092 0.163 0.160 0.859 1.081 0.855 0.851
Unconditional 0.048 0.132 0.034 0.063 0.160 0.139 0.366 0.408 0.856 0.852
   In-Context 0.046 0.122 0.062 0.058 0.129 0.129 0.364 0.419 0.857 0.852
@@
4.4 Qualitative Results.
Layout Case Study. We specifically select samples with unique layout styles (upward slant) in the test set. Figure 7 shows that although the unconditional layout generation method can produce structurally reasonable layouts, it completely neglects the layout style of the handwritten text line. In comparison, our method effectively mimics characteristics such as glyph spacing and the slant trend of handwriting, verifying the necessity of our layout generator.
Figure 7: Visualization of character bounding boxes generated by different methods, where green represents the ground truth.
Subjective Visualization Experiment. In Figure 8, we select three authors with distinctly different writing styles as imitation targets. We visualized the real samples alongside the generated counterparts. The left columns display the real samples and the right columns display the generated data. It can be observed that our model adeptly imitates the target samples, capturing both the overall layout style, such as spacing and slant, as well as the calligraphic style of individual characters.
We take one sample from the real samples and one from the synthesized samples to form a pair, as shown in Figure 9. We create ten similar pairs and ask 20 participants whether the upper and lower lines are written by the same person. More than ninety percent of the responses are ‘yes’, which demonstrates that it is difficult to distinguish between the real samples and the synthesized ones.
@@ 5 Conclusions and Limitations
In this paper, we tackle the stylized online handwritten Chinese text line generation task hierarchically, which is almost unexplored. Our model consists of a novel text line layout generator and a stylized diffusion-based character synthesizer. The text line layout generator can arrange glyph positions based on the text contents and writing habits of the given reference samples while the stylized diffusion-based character synthesizer can generate characters with specific categories and calligraphy styles. However, due to the fact that each character is generated independently, our approach encounters difficulties in mimicking styles with extensive cursive connections between characters. An end-to-end text line generation method might be required to address this issue. Furthermore, whether synthesized samples can serve as augmentation data to assist in achieving better results for the recognizer is an open question worth exploring. Last but not least, how to generate more complex handwritten data, such as mathematical formulas based on the hierarchical approach is also an intriguing topic. We consider these as our future work.
Figure 9: Examples of subjective experiment. Each pair consists of one row from the synthesized samples and the other row from its corresponding target real samples. Participants need to determine whether one pair was written by the same person.

### v3

@@ Abstract
Text plays a crucial role in the transmission of human civilization, and teaching machines to generate online handwritten text in various styles presents an interesting and significant challenge. However, most prior work has concentrated on generating individual Chinese fonts, leaving complete text line generation largely unexplored. In this paper, we identify that text lines can naturally be divided into two components: layout and glyphs. Based on this division, we designed a text line layout generator coupled with a diffusion-based stylized font synthesizer to address this challenge hierarchically. More concretely, the layout generator performs in-context-like learning based on the text content and the provided style references to generate positions for each glyph autoregressively. Meanwhile, the font synthesizer which consists of a character embedding dictionary, a multi-scale calligraphy style encoder and a 1D U-Net based diffusion denoiser will generate each font on its position while imitating the calligraphy style extracted from the given style references. Qualitative and quantitative experiments on the CASIA-OLHWDB demonstrate that our method is capable of generating structurally correct and indistinguishable imitation samples. Our source code will be publicly available at: https://github.com/singularityrms/OLHWG .
@@ 1 Introduction
In this work, we aim to generate stylized and coherent online Chinese handwritten text lines based on specified text contents and style reference samples. A handwritten line of text can be viewed as composed of individual characters and their arrangement. Based on this observation, we propose a hierarchical method that disentangles the layout generation from the writing of individual characters during the generation of full lines. We design a novel text line layout generator to arrange positions for each character based on their categories and given handwriting style references. For font generation, we construct a 1D U-Net (Ronneberger et al., 2015) network and design a multi-scale feature Wang et al. (2023) fusion module to fully mimic the calligraphy characteristics of the given reference sample. We utilize the publicly CASIA-OLHWDB dataset (Liu et al., 2011) for training and evaluating our method, showcasing its effectiveness through a combination of quantitative and qualitative experiments. Figure 1 shows several visualizations of our generated samples.
Our contributions can be summarized as: 1) We propose a hierarchical approach to solve the unexplored online handwritten Chinese text line generation task. 2) We introduce a simple but effective layout generator that can generate character positions based on the text contents and writing style through in-context-like learning. 3) We construct a 1D U-Net-based network for font generation with a multi-scale contrastive-learning-based style encoder to improve the ability of calligraphy style imitation.
@@ 4.2 Font Generation
Baselines and Results. There are mainly two types of online handwritten Chinese character generation approaches: the first type comprises purely data-driven methods that learn the structure of characters and calligraphy style only from the training set. The second type requires a standard printed font image as an auxiliary input, and train models to transform the standard fonts into characters with a specific writing style. The advantage of the second method is that it is easier to generate correct structures, but it cannot generate characters without corresponding standard glyphs and the computational cost will be heavier. As shown in Table 1, being a method of the first type, our model achieves state-of-the-art performance in all pure data-driven approaches and is comparable to state-of-the-art style transfer-based methods. The qualitative comparison experiment is shown in Appendix A.2.5.
@@ 4.3.2 Quantitative Results
   Gaussian 0.074 0.151 0.092 0.092 0.163 0.160 0.859 1.081 0.851 0.855
Unconditional 0.048 0.132 0.034 0.063 0.160 0.139 0.366 0.408 0.852 0.856
   In-Context 0.046 0.122 0.062 0.058 0.129 0.129 0.364 0.419 0.852 0.857
@@
4.3.3 Qualitative Results.
Layout Case Study. We specifically select samples with unique layout styles (upward slant) in the test set. Figure 7 shows that although the unconditional layout generation method can produce structurally reasonable layouts, it completely neglects the layout style of the handwritten text line. In comparison, our method effectively mimics characteristics such as glyph spacing and the slant trend of handwriting, verifying the necessity of our layout generator. Additional experimental results can be found in Appendix A.4.
Figure 7: Visualization of character bounding boxes generated by different methods, where green represents the ground truth. For our method, the bounding boxes of the first ten characters are treated as the context prefix.
Subjective Visualization Experiment. We construct a visual subjective test as shown in Figure 8. Fifteen pairs of similar samples are created, and 20 participants are asked to determine whether lines in one block were written by the same person. Over 85% of the responses are ”yes,” indicating that the distinction between real and synthesized samples is difficult to perceive. It can be observed that our model adeptly imitates the target samples, capturing both the overall layout style, such as spacing and slant, as well as the calligraphic style of individual characters.
Figure 8: Examples of subjective experiments. For each author block, one row contains real samples and two rows contain synthesized samples. Participants must judge whether the samples were written by the same person.
@@ 5 Conclusions and Limitations
In this paper, we tackle the stylized online handwritten Chinese text line generation task hierarchically, which is almost unexplored. Our model consists of a novel text line layout generator and a stylized diffusion-based character synthesizer. The text line layout generator can arrange glyph positions based on the text contents and writing habits of the given reference samples while the stylized diffusion-based character synthesizer can generate characters with specific categories and calligraphy styles. However, due to the fact that each character is generated independently, our approach encounters difficulties in mimicking styles with extensive cursive connections between characters. An end-to-end text line generation method might be required to address this issue, as discussed in Appendix A.3. Furthermore, whether synthesized samples can serve as augmentation data to assist in achieving better results for the recognizer is an open question worth exploring. Last but not least, how to generate more complex handwritten data, such as mathematical formulas based on the hierarchical approach is also an intriguing topic. We consider these as our future work.
Acknowledgment: This work is supported by the Natural Science Foundation of China (NSFC) Grant No. 62436009, No. 62276258, and U23B2029.
@@ A.2.4 Implementation of Evaluation Metrics
A.2.5 More visualization results
For character generation, we have added a visual comparison between our method and the state-of-the-art style transfer method SDT Dai et al. (2023). As shown in Figure 10 our method achieves comparable performance and effectively mimics the target calligraphy style.
Figure 10: The visual comparisons with the state-of-the-art method SDT.
In Figure 11, For full text line generation, we select three authors with distinctly different writing styles as imitation targets. We visualized the real samples alongside the generated counterparts. The left columns display the real samples and the right columns display the generated data.
A.3 The Possibility of Extending to an End-to-End Approach
We demonstrate the results of directly using our proposed 1D U-Net generation module by concatenating individual character embeddings into a full-line embedding for end-to-end generation. As shown in Figure 13, we specifically selected a handwriting style with connected strokes between characters, and the end-to-end model is capable of handling this. This also illustrates the potential of extending our proposed method to an end-to-end generation framework, demonstrating the generalization of our module.
Figure 13 shows the failure case, specifically, as the number of characters increases, structural errors tend to occur. In contrast, the decoupled layout and glyph method offers more stable training and inference. The next challenge will be to enhance the stability and ensure a more reliable generative process for end-to-end generation, which will be the focus of our future work.
Figure 12: The end-to-end generation paradigm can produce connective strokes between characters.
Figure 13: The instability of directly using the end-to-end method with longer character sequences, resulting in failure cases.
A.4 Effects of Layout Generator
We also tested using a portion of another text line written by the same author as a prefix to condition the layout generator for generation, with the results shown in Figure 14. It can be observed that the layouts generated by our proposed method still closely approximate the ground truth. This is because different samples from the same author typically share a consistent writing style. The key principle of our method is to ensure that the generated layouts align with the style of the input prefix, rather than requiring the prefix to be from the ground truth text lines. In contrast to previous method, our contributions and advantages lie in:
1. Compared to previous model-free method, our approach is capable of generating layouts that are closer to the real target. The reason is that model-free methods deeply rely on manually designed rules and features, which, especially in the case of Chinese text lines, are challenged by the varying sizes and shapes of different types of Chinese characters, as well as the significant differences in punctuation marks. This makes it difficult for earlier method to simultaneously capture the layout style. In contrast, our method, based on the next-token prediction paradigm, allows the model to naturally generate a layout where not only the positions of characters are reasonable, but the overall style of the layout is consistent with the given prefix, in a in-context manner.
2. In addition, in terms of generalization, the layout generator can consider both the horizontal and vertical relative positional relationships between characters, therefore is capable of handling simple 2D mathematical expressions, such as $e^{x}$ and $\frac{1}{2}$ without any modifications, while model-free methods are unable to achieve.
Figure 14: Layout generation visualization, where the prefix is taken from other text lines by the same author. The previous method is denoted as ’Gaussian’.
We have supplemented the quantitative experiments on layout generation presented in Table 5 as follows, where In-Context-gt refers to the case where we use the first 10 bounding boxes from the ground truth text line as the prefix, while In-Context-other refers to the case where we use 10 bounding boxes from another randomly picked text line written by the same author as the prefix. As can be seen, the performance is almost identical. The result suggests that as long as the layout style across different lines from the same author remains consistent, our method can perform effectively.
$\nabla_{1}$ $\nabla_{2}$ $\nabla_{3}$ $\nabla_{4}$ $\nabla_{5}$ $\nabla_{6}$ $\nabla_{7}$ $\nabla_{8}$
In-Context-gt 0.046 0.122 0.062 0.058 0.129 0.129 0.364 0.419
In-Context-other 0.047 0.124 0.057 0.058 0.130 0.132 0.363 0.422
Table 5: The performance quantitative evaluation results of the layout generator using other lines from the same author as prefixes.